# SPDFusion: An Infrared and Visible Image Fusion Network Based on a Non-Euclidean Representation of Riemannian Manifolds

### v2

@@ 3 The Proposed Method
In this section, we provide a detailed description of the SPDFusion network framework, the cross-modal SPD fusion strategy, and the SPD Manifold Attention Module, as well as an explanation of our training process and loss function.
@@ 3.1 The Network Framework
The proposed fusion network (SPDFusion) architecture is shown in Fig. 2(a), the source images are first encoded into overlapping patch vectors, and then the covariance of the encoded patch row vector matrix is computed via the ”Cross-Modal Fusion Strategy” to detect statistical relationships between different patches. Through covariance computation, we not only capture the feature relationships within a single modality but also use algorithms to highlight low-correlation yet information-rich features across modalities. (This part will be explained in detail in Section 3.2.) In this way, the statistical representations within and between modalities are not only based on the original image space but also incorporate deep statistical information across multiple modalities, providing a rich data basis for manifold learning. Furthermore, to ensure the non-Euclidean geometric structure of the image is preserved, our designed manifold learning module completes cross-modal attention interaction within the Riemannian space, and the learned weight coefficients are ”flattened” back to Euclidean space through logarithmic mapping. Specifically, we designed a Riemannian attention weighting module based on the principle of convolution operations, injecting manifold statistical relevance into the image. Finally, the weighted image features are subjected to deep feature extraction through a series of convolutional layers to obtain the final fusion result.
The source images $I_{ir}$ and $I_{vi}$ are the infrared image and visible light image, respectively. To acquire a global spatial information representation, we first segment each image into patches with overlapping regions. For each patch, we arrange all the pixels in a sequential order to form a row vector. For each image modality, a matrix of row vectors corresponding to the rearranged image block matrices can be obtained. Specifically, let $P_{ir}$ and $P_{vi}$ be the patch sets of the infrared and visible light images,with $p_{ir_{i}}$ being the $i$-th patch of $I_{ir}$, and $p_{vi_{i}}$ being the $i$-th patch of $I_{vi}$. For each $p_{ir_{i}}$ in $P_{ir}$, we can construct the row vector $x_{ir_{i}}$. And for each $p_{vi_{i}}$ in $P_{vi}$, we can also construct the row vector $x_{vi_{i}}$. Eventually, we obtain the matrix of row vectors $X_{ir}$ for the infrared image and $X_{vi}$ for the visible light image, which can be expressed as:
$X_{ir}=\begin{bmatrix}x_{ir_{1}}\\ x_{ir_{2}}\\ \vdots\\ x_{ir_{n}}\end{bmatrix},\quad X_{vi}=\begin{bmatrix}x_{vi_{1}}\\ x_{vi_{2}}\\ \vdots\\ x_{vi_{n}}\end{bmatrix}$ (5)
@@
subsequently, we computed the covariance matrix ”Covariance1” between the vectorized patches, which encapsulates the correlation representation information of different regions in the image, denoted as $S$. To ensure that the covariance matrix satisfies the basic properties of the SPD manifold, we perform an SVD on this symmetric semi-definite matrix, and add a small positive perturbation term $\epsilon$ to the eigenvalue matrix, to ensure that all the eigenvalues are greater than zero:
$Q=U(S+\epsilon*tr(S))V^{T}$ (6)
Figure 3: The framework of our fusion strategy.
where $\epsilon$ is a very small positive number, usually set as 0.001. $U$ and $V^{T}$ represent the orthogonal matrices obtained from the SVD of the original positive semi-definite matrix, respectively. $S$ represents the diagonal matrix obtained by singular value decomposition, and the resultant matrix $Q$ belongs to an SPD manifold $X_{0}$ in the Riemannian space. $X_{0}$ goes through the SPD manifold network to learn the deep statistical information of different modalities, and embeds the learned feature manifold into the Euclidean space for attention weighting:
$\phi_{S_{i}}=X_{k}\phi_{row},i\in\left\{ir,vi\right\}$ (7)
where $X_{k}$ represents the weight matrix, which can also be represented as ”Covariance2”. $\phi_{S}$ represents the final feature map obtained, and $\phi_{row}$ represents the matrix of row vectors.
@@
$F_{f}=conv_{3\times 3}\left(cat\left(\phi_{S_{ir}},\phi_{S_{vi}}\right)\right)$ (8)
Following the steps mentioned above, we obtain the final fusion result $F_{f}$.
@@ 3.2 Cross Modal SPD Fusion Strategy
The core of this strategy is to use a structured feature map stacking method, in which the feature maps of two different modalities are arranged vertically for interaction within the same dimensional space. Specifically, the two source images are firstly divided into patches and flattened into vectors as shown in Fig. 3. In our work, the patch size, the stride between blocks, and the overlapping block size are set to 16, 8, and 8, respectively. After the patches are flattened, we obtain the row vector matrices $X_{ir}$ and $X_{vi}$, corresponding to different modalities. The process of vertical arrangement can be described as constructing a new stacked matrix $X_{stacked}$, which merges $X_{ir}$ and $X_{vi}$ as follows:
$X_{stacked}=\begin{bmatrix}X_{ir}\\ X_{vi}\end{bmatrix}$ (9)
@@
where $X_{ir}\in\mathbb{R}^{m\times n}$, $X_{vi}\in\mathbb{R}^{p\times n}$ and $X_{stacked}\in\mathbb{R}^{\left(m+p\right)\times n}$. Here $m$ and $p$ denote the number of the row vectors for the two modalities, and $n$ represent the dimension of the row vector. For any two rows $X_{i}$ and $X_{j}$ $\left(i,j=1,2,\dots,m+p,i\neq j\right)$ in $X_{stacked}$, the covariance representation between the row vectors can be formulated as:
$Cov\left(X_{i},X_{j}\right)=\frac{1}{n-1}\sum_{k=1}^{n}\left(X_{ik}-\bar{X_{i}}\right)\left(X_{jk}-\bar{X_{j}}\right)^{T}$ (10)
@@
where $\bar{X_{i}}$ and $\bar{X_{j}}$ represent the average of all observed feature values for the $i$-th patch and the $j$-th patch, respectively.
@@
Then, a composite covariance matrix is constructed, serving as a vehicle for analyzing the collaborative variation trends between composite data points on the entire high-dimensional data manifold. The matrix is divided into four quadrant blocks:
$C_{X}=\begin{bmatrix}C_{X_{ir}}&C_{X_{ir}X_{vi}}\\ C_{X_{vi}X_{ir}}&C_{X_{vi}}\end{bmatrix}$ (11)
Among them,
$C_{X_{ir}}=\frac{1}{n-1}\sum_{k=1}^{n}\left(X_{ir_{k}}-\bar{X}_{ir}\right)\left(X_{ir_{k}}-\bar{X}_{ir}\right)^{T}$ (12)
$C_{X_{vi}}=\frac{1}{n-1}\sum_{k=1}^{n}\left(X_{vi_{k}}-\bar{X}_{vi}\right)\left(X_{vi_{k}}-\bar{X}_{vi}\right)^{T}$ (13)
@@
where $C_{X_{ir}}$ and $C_{X_{vi}}$ are the covariance matrices of the patches within $X_{ir}$ and $X_{vi}$ themselves, representing the self-correlation characteristics within a single modality. Simultaneously,
$C_{X_{ir}X_{vi}}=C_{X_{vi}X_{ir}}=\frac{1}{n-1}\sum_{k=1}^{n}\left(X_{ir_{k}}-\bar{X}_{ir}\right)\left(X_{vi_{k}}-\bar{X}_{vi}\right)^{T}$ (14)
where $C_{X_{ir}X_{vi}}$ and $C_{X_{vi}X_{ir}}$ represent the covariance matrices between $X_{ir}$ and $X_{vi}$, and $C_{X_{vi}X_{ir}}$ is the transpose of $C_{X_{ir}X_{vi}}$, representing the cross-modal covariance computation, which reveals the relationship between the two modalities.
The covariance matrix $C_{X}$ reveals the principle of mutual distribution of points on high-dimensional manifolds from a statistical perspective, explicitly expressing the dependence between different feature points within the SPD manifold. As a key component of cross-modal fusion, it breaks away from the traditional fusion strategy of simply adding through Euclidean distance, providing a new perspective for understanding the non-Euclidean interactions between various modal data.
Once these advanced relational representations are obtained, the model integrates them with the input features to derive the modality-associated enhanced features $\mathrm{IR}$ and $\mathrm{VI}$, thereby attracting the attention of the neural network. Furthermore, through gradient backpropagation on the manifold, the model continuously refines the parameters of the SPD matrix to facilitate cross-fusion.
Figure 4: Visualization of the training results on the manifold under different iteration counts. The first column represents the changes in the Riemannian attention weight matrix, while the second column shows the intermediate results of the fusion.
@@ 3.4 SPD Manifold Attention Module
Following the conventional spatial attention mechanism, we introduce a symmetric positive definite attention module (SPDAM), which hinges on exploiting global semantic associations intrinsic to Riemannian manifolds for enhanced attention modeling. The core of SPDAM lies in its adept usage of a covariance matrix that not only encapsulates symmetry and positive definiteness, but also facilitates the reallocation of manifold attention weights in the Euclidean space.
Central to our method is the concept that points within the SPD manifold, through the innovative deployment of SPDAM, are effectively ”flattened” or linearized, enabling their approximation within a Euclidean framework. Mathematically, this transformation can be conceptualized as mapping points $p$ on the manifold $M$ through a function $f:M\to\mathbb{R}^{d}$ that linearizes these points within the Euclidean space $\mathbb{R}^{d}$.
Moreover, each pixel in SPD features represents the statistical correlation between feature dimensions, which is a global relationship and can be used as a weight to weight the original image features. In addition, the learning of weight parameters in this article is carried out in the intrinsic space of the data, so the learned weights may be more reasonable. The ablation experiment in the following text also supports our speculation.
As illustrated on the Fig. 2(c), the element at row $i$ and column $j$ in the symmetric matrix generated by our SPD network can be viewed as the statistical correlation between the $i$-th patch and the $j$-th patch. By multiplying the learned weight matrix with the blocked row vector matrix, the weighted matrix reflects the aggregation of covariance weights of the $j$-th feature (i.e., the $j$-th pixel position) in the $i$-th patch with the same feature in all patches across the image. More specifically, it captures the information aggregation achieved by weighting local features in the global context through the covariance relationships.
@@ 3.5 Loss Function
Our SPDFusion is constrained by int loss $L_{int}$, grad loss $L_{grad}$ , structure similarity loss $L_{ssim}$ and covariance loss $L_{cov}$. This is done to enable the network to further enhance semantic information through the correlation between deep features, while learning the color brightness information and texture details in the image. The total loss function $L_{total}$ is defined as follows:
$L_{total}=L_{int}+\alpha L_{grad}+\beta L_{ssim}+\gamma L_{cov}$ (15)
$L_{int}$ calculates the difference between the generated image and the target image at the pixel level. It reflects the difference in intensity between images, and is defined as:
$L_{int}=\frac{1}{HW}\parallel I_{f}-max(I_{ir},I_{vis})\parallel_{1}$ (16)
@@
$L_{grad}=\frac{1}{HW}\parallel\left|\nabla I_{f}\right|-max(\left|\nabla I_{ir}\right|,\left|\nabla I_{vis}\right|)\parallel_{1}$ (17)
where $\nabla$ is the Sobel operator, and ||$\cdot$|| ⋅ | represents the absolute value operation.
@@
$L_{ssim}=\left(1-ssim\left(I_{f},I_{vis}\right)\right)+\left(1-ssim\left(I_{f},I_{ir}\right)\right)$ (18)
@@
Furthermore, to restrain the highly correlated features of the image, we employ VGG-16 [34], trained on ImageNet, to extract features. The third and fourth layers of the VGG network are selected to devise the loss function, which allows us to capture a more abstract characterization of the correlations. $L_{cov}$ is defined as follows:
$L_{cov}=\sum_{k=3}^{4}||Cov(\phi(I_{f})^{k})-Cov(\phi(I_{ir})^{k})||_{1}$ (19)
Here, $Cov\left(\cdot\right)$ denotes the covariance matrix of the feature map. Its diagonal elements represent the auto-correlation of each feature appearing in the image, while the off-diagonal elements indicate the correlations between different features.
TNO
MSRS
Figure 5: The examples of two datasets: TNO and MSRS.
(a) Infrared image
(b) Visible image
(c) DenseFuse
(d) FusionGAN
(e) GANMcC
(f) RFN-Nest
(g) DeFusion
(h) FLFuse
(i) SemLA
(j) LRRNet
(k) DDFM
(l) SPDFusion
(a) Infrared image
(b) Visible image
(c) DenseFuse
(d) FusionGAN
(e) GANMcC
(f) RFN-Nest
(g) DeFusion
(h) FLFuse
(i) SemLA
(j) LRRNet
(k) DDFM
(l) SPDFusion

### v3

@@ 3 The Proposed Method
In this section, we provide a detailed description of the SMLNet network framework, the cross-modal SPD fusion strategy, the manifold learning network and the SPD attention module (SPDAM), as well as an explanation of our training process and loss function.
@@ 3.1 The Network Framework
The proposed fusion network (SMLNet) architecture is shown in Fig. 2(a), the source images are first encoded into overlapping patch vectors, and then the covariance of the encoded patch row vector matrix is computed via the “Cross-Modal Fusion Strategy” to detect statistical relationships between different patches. Through covariance computation, we not only capture the feature relationships within a single modality but also use algorithms to highlight low-correlation yet information-rich features across modalities. (This part will be explained in detail in Section 3.2) In this way, the statistical representations within and between modalities are not only based on the original image space but also incorporate deep statistical information across multiple modalities, providing a rich data basis for manifold learning. Furthermore, to ensure the non-Euclidean geometric structure of the image, our manifold learning module completes cross-modal attention interaction within the Riemannian space, and the learned weight coefficients are “flattened” back to Euclidean space through logarithmic mapping. Specifically, we designed a Riemannian attention weighting module based on the principle of convolution operations, injecting manifold statistical relevance into the image. Finally, the weighted image features are subjected to deep feature extraction through a series of convolutional layers to obtain the final fusion result.
The source images $I_{\text{ir}}$ and $I_{\text{vi}}$ are the infrared image and visible light image, respectively. To acquire a global spatial information representation, we first segment each image into patches with overlapping regions. For each patch, we arrange all the pixels in a sequential order to form a row vector. For each image modality, a matrix of row vectors corresponding to the rearranged image block matrices can be obtained. Specifically, let $P_{\text{ir}}$ and $P_{\text{vi}}$ be the patch sets of the infrared and visible light images,with $p_{\text{ir}_{i}}$ being the $i$-th patch of $I_{\text{ir}}$, and $p_{\text{vi}_{i}}$ being the $i$-th patch of $I_{\text{vi}}$. For each $p_{\text{ir}_{i}}$ in $P_{\text{ir}}$, we can construct the row vector $x_{\text{ir}_{i}}$. And for each $p_{\text{vi}_{i}}$ in $P_{\text{vi}}$, we can also construct the row vector $x_{\text{vi}_{i}}$. Eventually, we obtain the matrix of row vectors $\mathbf{M}_{\text{ir}}$ for the infrared image and $\mathbf{M}_{\text{vi}}$ for the visible light image, which can be expressed as:
$\mathbf{M}_{\text{ir}}=\begin{bmatrix}x_{\text{ir}_{1}}\\ x_{\text{ir}_{2}}\\ \vdots\\ x_{\text{ir}_{n}}\end{bmatrix},\quad\mathbf{M}_{\text{vi}}=\begin{bmatrix}x_{\text{vi}_{1}}\\ x_{\text{vi}_{2}}\\ \vdots\\ x_{\text{vi}_{n}}\end{bmatrix}$, (5)
@@
Subsequently, we computed the covariance matrix “Covariance1” between the vectorized patches, which encapsulates the correlation representation information of different regions in the image. To ensure that the covariance matrix satisfies the basic properties of the SPD manifold, we perform an SVD on this symmetric semi-definite matrix, and add a small positive perturbation term $\epsilon$ to the eigenvalue matrix, to ensure that all the eigenvalues are greater than zero:
$Q=U(S+\epsilon*tr(S))V^{T},$ (6)
where $\epsilon$ is a very small positive number, usually set as 0.001. $U$ and $V^{T}$ represent the orthogonal matrices obtained from the SVD of the original positive semi-definite matrix, respectively. $S$ represents the diagonal matrix obtained by singular value decomposition, and the resultant matrix $Q$ belongs to an SPD manifold $X_{0}$ in the Riemannian space. Then, $X_{0}$ goes through the SPD manifold network to learn the deep statistical information of different modalities, and embeds the learned feature manifold into the Euclidean space for attention weighting:
$F_{\text{mod}}=\text{MatMul}(X_{k},\phi_{\text{row}}),\quad\text{mod}\in\{\text{ir},\text{vi}\}$ (7)
where $X_{k}$ represents the weight matrix, which can also be represented as “Covariance2”. $F_{\text{mod}}$ represents the final feature map obtained, and $\phi_{\text{row}}$ represents the matrix of row vectors.
Figure 3: The of our fusion strategy. The vectorized image patches are modeled as covariance matrices that represent spatial relationship attributes. Attention weights are adaptively allocated through manifold learning to promote information fusion within a single modality and between modalities.
Figure 4: Visualization of training results at different iterations. The first row represents the source image. The second and fourth rows show intermediate fusion results, while the third and fifth rows display the evolution of Riemannian attention weight matrices. Initially, the weights concentrate on intra-modality (diagonal regions), yielding blurry fused outputs. As training progresses, the weights gradually spread to inter-modality (off-diagonal regions), demonstrating the network’s ability to integrate cross-modal information. This evolution produces more structurally coherent fused images.
@@
$I_{\text{f}}=conv_{3\times 3}\left(cat\left(F_{\text{ir}},F_{\text{vi}}\right)\right),$ (8)
where $I_{\text{f}}$ denotes the final fusion result obtained through the above steps.
@@ 3.2 Cross Modal SPD Fusion Strategy
The core of this strategy is to use a structured feature map stacking method, in which the feature maps of two different modalities are arranged vertically for interaction within the same dimensional space. Specifically, the two source images are firstly divided into patches and flattened into vectors as shown in Fig. 3. In our work, the patch size, the stride between blocks, and the overlapping block size are set to 16, 8, and 8, respectively. After the patches are flattened, we obtain the row vector matrices $\mathbf{M}_{\text{ir}}$ and $\mathbf{M}_{\text{vi}}$, corresponding to different modalities. The process of vertical arrangement can be described as constructing a new stacked matrix $\mathbf{M}_{\text{stacked}}$, which merges $\mathbf{M}_{\text{ir}}$ and $\mathbf{M}_{\text{vi}}$ as follows:
$\mathbf{M}_{\text{stacked}}=\begin{bmatrix}\mathbf{M}_{\text{ir}}\\ \mathbf{M}_{\text{vi}}\end{bmatrix}$, (9)
@@
where $\mathbf{M}_{\text{ir}}\in\mathbb{R}^{m\times n}$, $\mathbf{M}_{\text{vi}}\in\mathbb{R}^{p\times n}$ and $\mathbf{M}_{\text{stacked}}\in\mathbb{R}^{\left(m+p\right)\times n}$. Here $m$ and $p$ denote the number of the row vectors for the two modalities, and $n$ represent the dimension of the row vector. For any two rows $\mathbf{M}_{i}$ and $\mathbf{M}_{j}$ $\left(i,j=1,2,\dots,m+p,i\neq j\right)$ in $\mathbf{M}_{\text{stacked}}$, the covariance representation between the row vectors can be formulated as:
$Cov\left(\mathbf{M}_{i},\mathbf{M}_{j}\right)=\frac{1}{n-1}\sum_{k=1}^{n}\left(\mathbf{M}_{ik}-\bar{\mathbf{M}}_{i}\right)\left(\mathbf{M}_{jk}-\bar{\mathbf{M}}_{j}\right)^{T},$ (10)
@@
where $\bar{\mathbf{M}}_{i}$ and $\bar{\mathbf{M}}_{j}$ represent the average of all observed feature values for the $i$-th patch and the $j$-th patch, respectively.
@@
Then, a composite covariance matrix is constructed to analyze the collaborative variation trends between composite data points on the high-dimensional data manifold. The matrix is divided into four quadrant blocks:
$\mathbf{C}_{\mathbf{M}}=\begin{bmatrix}\mathbf{C}_{\mathbf{M}_{\text{ir}}}&\mathbf{C}_{\mathbf{M}_{\text{ir}}\mathbf{M}_{\text{vi}}}\\ \mathbf{C}_{\mathbf{M}_{\text{vi}}\mathbf{M}_{\text{ir}}}&\mathbf{C}_{\mathbf{M}_{\text{vi}}}\end{bmatrix}$, (11)
where
$\mathbf{C}_{\mathbf{M}_{\text{ir}}}=\frac{1}{n-1}\sum_{k=1}^{n}\left(\mathbf{M}_{\text{ir}_{k}}-\bar{\mathbf{M}}_{\text{ir}}\right)\left(\mathbf{M}_{\text{ir}_{k}}-\bar{\mathbf{M}}_{\text{ir}}\right)^{T},$ (12)
and
$\mathbf{C}_{\mathbf{M}_{\text{vi}}}=\frac{1}{n-1}\sum_{k=1}^{n}\left(\mathbf{M}_{\text{vi}_{k}}-\bar{\mathbf{M}}_{\text{vi}}\right)\left(\mathbf{M}_{\text{vi}_{k}}-\bar{\mathbf{M}}_{\text{vi}}\right)^{T}.$ (13)
@@
Here, $\mathbf{C}_{\mathbf{M}_{\text{ir}}}$ and $\mathbf{C}_{\mathbf{M}_{\text{vi}}}$ are the covariance matrices of the patches within $\mathbf{M}_{\text{ir}}$ and $\mathbf{M}_{\text{vi}}$, respectively, representing the self-correlation characteristics of each modality. The cross-modal covariance is given by
$\begin{split}\mathbf{C}_{\mathbf{M}_{\text{ir}}\mathbf{M}_{\text{vi}}}&=\mathbf{C}_{\mathbf{M}_{\text{vi}}\mathbf{M}_{\text{ir}}}\\ &=\frac{1}{n-1}\sum_{k=1}^{n}\left(\mathbf{M}_{\text{ir}_{k}}-\bar{\mathbf{M}}_{\text{ir}}\right)\left(\mathbf{M}_{\text{vi}_{k}}-\bar{\mathbf{M}}_{\text{vi}}\right)^{T},\end{split}$ (14)
where $\mathbf{C}_{\mathbf{M}_{\text{ir}}\mathbf{M}_{\text{vi}}}$ and $\mathbf{C}_{\mathbf{M}_{\text{vi}}\mathbf{M}_{\text{ir}}}$ capture the relationship between the two modalities, with $\mathbf{C}_{\mathbf{M}_{\text{vi}}\mathbf{M}_{\text{ir}}}$ being the transpose of $\mathbf{C}_{\mathbf{M}_{\text{ir}}\mathbf{M}_{\text{vi}}}$.
This Riemannian representation fundamentally differs from Euclidean feature addition, as trivial summation fails to capture higher-order statistical interactions. In contrast, $\mathbf{C}_{\mathbf{M}}$ preserves the complex nonlinear dependencies between modalities through its tensor structure. During backpropagation, the gradients constrained by the manifold geometry adaptively refine $\mathbf{C}_{\mathbf{M}}$, optimizing the intrinsic connections between modalities. This process yields enhanced features $\mathrm{IR}$ and $\mathrm{VI}$ with geometrically consistent fusion results.
@@ 3.3 Manifold Learning Network
TNO
MSRS
RoadScene
Figure 5: The examples of three datasets: TNO, MSRS and RoadScene.
Figure 8: Infrared and visible image fusion experiment on RoadScene dataset.
@@ 3.4 SPD Manifold Attention Module
Following conventional spatial attention, we introduce the Symmetric Positive Definite Attention Module (SPDAM) to exploit global semantic associations within Riemannian manifolds. Central to SPDAM is its covariance matrix, which inherently encapsulates symmetry and positive definiteness while enabling manifold attention redistribution in Euclidean space. This mechanism effectively “flattens” points on the SPD manifold through a mapping function $f:\mathcal{M}\rightarrow\mathbb{R}^{d}$, linearly approximating manifold structures within Euclidean coordinates.
Crucially, as shown in Fig. 2(c), each element in our SPD feature matrix quantifies statistical correlations between the $i$-th and $j$-th patches. Based on this, we integrate cross-modal statistical correlation weights by multiplying the learned weight matrix with blocked row vectors. This operation specifically weights the $j$-th pixel position (corresponding locations across all patches) into the $i$-th patch’s features, thereby restructuring local features within a global context. Furthermore, we learn weight parameters directly in the manifold space, yielding more principled weighting schemes. Later ablation studies empirically validate this design advantage.
@@ 3.5 Loss Function
Our SMLNet is constrained by int loss $L_{\text{int}}$, grad loss $L_{\text{grad}}$ , structure similarity loss $L_{\text{ssim}}$ and covariance loss $L_{\text{cov}}$. This is done to enable the network to further enhance semantic information through the correlation between deep features, while learning the color brightness information and texture details in the image. The total loss function $L_{\text{total}}$ is defined as follows:
$L_{\text{total}}=L_{\text{int}}+\alpha L_{\text{grad}}+\beta L_{\text{ssim}}+\gamma L_{\text{cov}},$ (15)
where $L_{\text{int}}$ calculates the difference between the generated image and the target image at the pixel level. It reflects the difference in intensity between images, and is defined as:
$L_{\text{int}}=\frac{1}{HW}\parallel I_{\text{f}}-max(I_{\text{ir}},I_{\text{vi}})\parallel_{1},$ (16)
@@
$L_{\text{grad}}=\frac{1}{HW}\parallel\left|\nabla I_{\text{f}}\right|-max(\left|\nabla I_{\text{ir}}\right|,\left|\nabla I_{\text{vi}}\right|)\parallel_{1},$ (17)
where $\nabla$ is the Sobel operator, and $|\cdot|$ represents the absolute value operation.
@@
$L_{\text{ssim}}=\left(1-ssim\left(I_{\text{f}},I_{\text{vi}}\right)\right)+\left(1-ssim\left(I_{\text{f}},I_{\text{ir}}\right)\right),$ (18)
@@
Furthermore, to restrain the highly correlated features of the image, we employ VGG-16  (Simonyan and Zisserman, 2014), trained on ImageNet, to extract features. The third and fourth layers of the VGG network are selected to devise the loss function, which allows us to capture a more abstract characterization of the correlations. $L_{\text{cov}}$ is defined as follows:
[…]
where $Cov\left(\cdot\right)$ denotes the covariance matrix of the feature map. Its diagonal elements represent the auto-correlation of each feature appearing in the image, while the off-diagonal elements indicate the correlations between different features. The $k$ denotes the network depth.
Table 1: Quantitative Experiments on the TNO Dataset. We highlight the top three metrics using bolditalic, bold and italic fonts.
Methods Year EN MI SD SF VIF AG $Q^{AB/F}$
DenseFuse (Li and Wu, 2018) 2018 6.380 2.227 25.259 6.570 0.574 2.569 0.349
RFN-Nest (Li et al., 2021a) 2021 6.971 2.126 37.090 5.931 0.550 2.691 0.333
DeFusion (Liang et al., 2022) 2022 6.607 2.627 30.868 6.461 0.559 2.624 0.365
SemLA (Xie et al., 2023) 2023 6.879 2.051 37.890 8.528 0.470 2.892 0.265
DDFM (Zhao et al., 2023b) 2023 6.854 2.231 34.418 8.629 0.631 3.405 0.434
LRRNet (Li et al., 2023) 2023 6.991 2.518 40.984 9.608 0.548 3.790 0.352
IRFS (Wang et al., 2023b) 2023 6.651 2.150 31.784 8.939 0.587 3.242 0.397
CrossFuse (Li and Wu, 2024) 2024 6.908 2.952 39.976 9.951 0.716 3.733 0.425
VDMUFusion (Shi et al., 2024) 2024 6.358 2.076 24.980 4.832 0.502 2.048 0.249
DCINN (Wang et al., 2024) 2024 6.772 2.143 34.977 9.626 0.364 3.837 0.290
DDBFusion (Zhang et al., 2025) 2025 6.620 1.935 37.343 10.585 0.511 4.143 0.350
FusionBooster (Cheng et al., 2025b) 2025 6.516 2.455 32.863 10.898 0.480 3.455 0.319
GIFNet (Cheng et al., 2025a) 2025 6.974 1.990 41.429 13.742 0.501 5.132 0.347
SMLNet Ours 7.068 3.894 43.209 11.098 0.832 4.301 0.516
Table 2: Quantitave Experiments on the MSRS Dataset. The top three metrics are displayed in bolditalic, bold and italic fonts, respectively.
Methods Years EN MI SD SF VIF AG $Q^{AB/F}$
DenseFuse (Li and Wu, 2018) 2018 5.931 2.666 23.550 6.020 0.692 2.053 0.368
RFN-Nest (Li et al., 2021a) 2021 6.196 2.460 29.078 6.163 0.656 2.115 0.390
DeFusion (Liang et al., 2022) 2022 6.383 2.990 35.429 8.146 0.730 2.654 0.507
SemLA (Xie et al., 2023) 2023 6.423 2.461 33.122 6.351 0.617 2.257 0.290
DDFM (Zhao et al., 2023b) 2023 6.175 2.735 28.925 7.388 0.743 2.522 0.474
LRRNet (Li et al., 2023) 2023 6.192 2.928 31.758 8.473 0.541 2.651 0.454
IRFS (Wang et al., 2023b) 2023 6.603 2.159 35.868 9.888 0.735 3.155 0.477
CrossFuse (Li and Wu, 2024) 2024 6.494 3.132 36.447 9.661 0.839 3.019 0.561
VDMUFusion (Shi et al., 2024) 2024 5.948 2.331 23.600 6.490 0.633 2.242 0.344
DCINN (Wang et al., 2024) 2024 6.044 3.054 40.288 10.783 0.721 3.473 0.509
DDBFusion (Zhang et al., 2025) 2025 5.971 2.210 28.422 8.551 0.628 2.791 0.385
FusionBooster (Cheng et al., 2025b) 2025 6.316 2.075 30.215 8.801 0.639 3.051 0.422
GIFNet (Cheng et al., 2025a) 2025 5.940 1.971 32.900 12.705 0.582 3.367 0.416
SMLNet Ours 6.478 3.903 41.571 9.965 0.855 3.063 0.566
Table 3: Quantitave Experiments on the RoadScene Dataset. For the best three metrics, we use bolditalic, bold and italic fonts to distinguish them.
Methods Years EN MI SD SF VIF AG $Q^{AB/F}$
DenseFuse (Li and Wu, 2018) 2018 6.765 2.939 31.384 5.211 0.610 2.226 0.389
RFN-Nest (Li et al., 2021a) 2021 7.253 2.877 44.926 6.641 0.717 2.733 0.375
DeFusion (Liang et al., 2022) 2022 6.906 3.092 34.790 5.437 0.601 2.323 0.412
SemLA (Xie et al., 2023) 2023 6.954 2.856 38.566 8.119 0.564 2.515 0.271
DDFM (Zhao et al., 2023b) 2023 7.242 2.952 43.262 7.223 0.676 3.080 0.483
LRRNet (Li et al., 2023) 2023 7.145 2.866 44.110 8.196 0.578 3.299 0.382
IRFS (Wang et al., 2023b) 2023 7.010 2.999 37.631 6.385 0.649 2.695 0.476
CrossFuse (Li and Wu, 2024) 2024 7.189 3.520 47.273 7.578 0.663 3.055 0.366
VDMUFusion (Shi et al., 2024) 2024 6.798 2.845 32.242 5.665 0.576 2.358 0.375
DCINN (Wang et al., 2024) 2024 6.940 3.313 36.681 7.526 0.706 3.046 0.480
DDBFusion (Zhang et al., 2025) 2025 6.880 2.723 44.499 8.668 0.633 3.597 0.438
FusionBooster (Cheng et al., 2025b) 2025 6.524 3.127 29.351 6.098 0.495 2.519 0.347
GIFNet (Cheng et al., 2025a) 2025 7.241 2.659 46.338 11.460 0.576 4.293 0.413
SMLNet Ours 7.270 5.055 56.767 8.549 1.104 3.139 0.415